# Transfer Learning of Surrogate Models: Integrating Domain Warping and Affine Transformations

### v1

@@ Abstract.
conference: Genetic and Evolutionary Computation Conference (GECCO ’25); July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25), July 14–18, 2025, Malaga, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Transfer learningccs: Theory of computation Gaussian processesccs: Theory of computation Continuous optimization
@@ 1. Introduction
Surrogate modeling (Forrester et al., 2008; Forrester and Keane, 2009; Bhosekar and Ierapetritou, 2018; Tong et al., 2021) is extensively used to replace expensive simulators for reducing computational costs, for instance in automobile industry (Kiani and Yildiz, 2016; Fang et al., 2017; Qiu et al., 2018; Thomaser et al., 2023). Machine learning models are commonly used as surrogates, e.g., Gaussian process regression (GPR) (Satria Palar et al., 2020; Rajaram et al., 2020; Pourmohamad, 2021) and random forest (Wang and Jin, 2020; Antoniadis et al., 2021). Training a surrogate model on a new problem often requires a large number of training samples, in particular when there are many independent variables. Acquiring this data usually implies running expensive simulations or real-world experiments. Therefore, we wish to avoid the cost of acquiring large data sets to build surrogates on a new problem instance. Transfer learning (Pan and Yang, 2010; Weiss et al., 2016; Zhuang et al., 2021) can be used to tackle this issue: with a tiny transfer data set sampled on a new problem (the target), we can learn to tweak an accurate surrogate trained on an old problem (the source), provided certain symmetry/invariances between problems.
Covariance shift (Sugiyama et al., 2007; Shimodaira, 2000; Pathak et al., 2022) is an important type of symmetry, which says that for a regression task to approximate the source function $f^{\text{S}}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$, a target function $f^{\text{T}}$ can be obtained from $f^{\text{S}}$ by transforming the domain thereof. Namely, there exists a bijection $g\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that $f^{\text{T}}=f^{\text{S}}\circ g$. As for the surrogate modeling, the covariance shift implies the predictive distribution $P(y|\mathbf{x})$ remains unchanged between the source and the target while $P(\mathbf{x})$ differs. Previous studies (Pan et al., 2024, 2025) have investigated a special case: $g$ being an affine transformation, where the transformation is learned by minimizing a loss function on a tiny transfer data evaluated on the target.
@@
Figure 1. On the two-dimensional F7 StepEllipsoid function, we show, from left to right, the contour lines of the source function ($f^{\text{S}}$), the target ($f^{\text{T}}$), the original GPR ($\hat{f}^{\mkern 1.0mu\text{S}}$) trained to approximate $f^{\text{S}}$, the GPR model trained from scratch with 40 points sampled on $f^{\text{T}}$, and the original GPR model transferred with these 40 points. $f^{\text{T}}$ is created from $f^{\text{S}}$ by transforming the domain thereof with affine warping. We show the transfer effect with 80 data points in the last two subplots.
@@
However, the affine transformation $g$ might be too restrictive to model complex real-world scenarios, e.g., non-linearity is necessary in the automobile industry problems (Pan et al., 2025). Hence, we propose implementing a non-linear $g$ function with a beta cumulative distribution function (CDF) (Snoek et al., 2014). We showcase the preliminary results of our method in Figure 1, where we train a Gaussian process regression (GPR) model train on function F7 from the BBOB benchmark suite. Comparing the transferred model to the one trained from scratch on the target function (fourth and fifth subplots), we see the transferred model is more accurate and resembles the contour lines of the target function. Our contributions are:
@@
The paper is organized as follows: Sec. 2 briefly reviews the related works. In Sec. 3, we describe our methodology, followed by details of the experimental setup and synthetic and real-world transfer learning problems (Sec. 4). In Sec. 5, we provide a comprehensive analysis and discussion of the experimental results. Sec. 6 summarizes the key empirical findings and points out the next steps.
@@ 2. Related Works
Saida and Nishio (2023) investigated GPR surrogate model transfer for structural reliability under uncertainties by augmenting the feature space (Daumé, 2009). Zhang et al. (2023) proposed a novel transfer learning strategy that employs a geodesic flow kernel and knee point-based manifold learning to refine Gaussian process models using high-quality knee solutions from previous tasks, thereby enriching training data and boosting solution precision. In multi-task learning, Cao et al. (2010) developed the Adaptive Transfer Learning algorithm (AT-GP) using a semi-parametric transfer kernel. The Transfer Bayesian Committee Machine (Tr-BCM) (Da et al., 2019) introduced a scalable transfer learning approach by aggregating predictions from lightweight local experts, relaxing assumptions of uniform similarity between tasks. Recent advances include Papez and Quinn (2022), which proposed a probabilistic predictor for global source-target interactions, and Wei et al. (2023), which developed an interpretable multi-source transfer kernel for improved cross-task performance.
@@
Snoek et al. (2014) propose warping the independent variables with beta CDF to realize non-stationary kernels, where the unknown shape parameters of the beta distribution are inferred with Bayesian estimation (using log-normal priors). Their posterior predictive distribution is obtained by marginalizing the shape parameters. In contrast, in this work, we use the beta CDF to model the non-linear relation between the domain of the source and the target function. Also, we learn the unknown shape parameters with a loss-minimization approach.
Cowen-Rivers et al. (2022) introduced Kumaraswamy input warping, offering a more computationally efficient alternative to the beta CDF with similar flexibility. Du et al. (2017) presented a Hypothesis Transfer Learning framework linking domains via transformation functions, and Zhu et al. (2023) proposed a nonlinear transformation method to align the marginal probability distributions without prior data knowledge.
@@ 3. Learning Affine Warping to transfer domains
where $B(\alpha_{i},\beta_{i})$ is the beta function. The beta CDF parameterization for each dimension has two unknown parameters $\alpha_{i}>0,\beta_{i}>0$. We illustrate the effect of input warping for a 2D sphere function in Fig. 2, where each axis is warping non-linearly with different shape parameters.
Figure 2. 2D input warping. The coordinate system is transformed from left to right by beta CDFs with shape parameter $\alpha=1.0558,\beta=1.9339$ for $x$-axis and $\alpha=0.8655,\beta=1.8148$ for $y$-axis. We show the contour lines of a sphere function on the left and its warped version on the right.
@@ 3.2. Transfer non-differentiable surrogates
We also wish to apply our methodology to non-differentiable models like random forests. We propose to use the Covariance matrix adaptation evolution strategy (CMA-ES) (Emmerich et al., 2018; Hansen, 2006, 2016) to tune the parameters. CMA-ES can be applied directly to the search space of the translation parameter $\mathbf{v}$ and the beta CDF parameters $\mathbf{\alpha}$ and $\mathbf{\beta}$, which are Euclidean. However, special treatment is needed for $\mathbf{W}$, which lives in a smooth manifold $\operatorname{SO}\!\left(d\right)$. To solve this issue, we consider the Lie group representation $\mathfrak{so}(d)=\{\mathbf{A}\in\mathbb{R}^{d\times d}\colon\mathbf{A}^{\top}=-\mathbf{A}\}$, which is a flat space (with dimension $d(d-1)/2$), and optimize this representation with CMA-ES. A rotation matrix $\mathbf{W}$ can be recovered from its representation $\mathbf{A}$ with the exponential map, i.e., $\mathbf{W}=\operatorname{Exp}(\mathbf{A})$. For each search point $\mathbf{z}\in\mathbb{R}^{d(d-1)/2}$, we have to transform it into a $d\times d$ antisymmetric matrix to preserve the structure of $\mathfrak{so}(d)$: the components in $\mathbf{z}$ are sequentially assigned to the upper triangular of $\mathbf{A}$ (diagonal entries are zero) row by row. The negative value of the transposition of the upper triangular then fills the lower triangular entries.
4. Experimental Settings
Figure 3. On 2D BBOB functions, we compare the transferred GPR models with those trained from scratch on the transfer dataset. Each cell displays the percentage difference in average SMAPE (%) for a combination of BBOB functions, sample size, and beta CDF shape. Positive values (shown in red) indicate superior performance of the transferred model and vice versa.
@@ 4. Experimental Settings
We first evaluate our method on the Black-Box Optimization Benchmarking (BBOB) (Hansen et al., 2009a, 2020, b) suite, which consists of 24 continuous, single-objective problems. The BBOB suite has been widely used as a regression benchmark (Singh and Dwivedi, 2018; Tian et al., 2020; Chen et al., 2022; Yang and Affenzeller, 2023) as it reflects real-world regression difficulties. To create synthetic transfer learning problems out of BBOB, we take the first problem instance of each BBOB function as the source $f^{\text{S}}$, and construct the target $f^{\text{T}}$ by applying a beta CDF transformation, followed by random rotation and translation transformations, to the base function. The shape parameters are sampled from log-normal distributions, i.e., $\log\alpha_{i}\sim\mathcal{N}(\mu_{i}^{\alpha},\sigma^{\alpha}_{i}),\log\beta_{i}\sim\mathcal{N}(\mu_{i}^{\beta},\sigma_{i}^{\beta})$. Inspired by (Snoek et al., 2014), we choose different priors to realize distinct shapes of the beta CDF:
@@
This dataset evaluates the performance of optimization algorithms in the context of automotive engineering, with a specific focus on minimizing braking distances. It includes five distinct vehicle configurations, each characterized by unique combinations of tire performance and vehicle load conditions. The search space consists of two Anti-lock Braking System (ABS) control parameters, $x_{1}$ and $x_{2}$, which span a total of 10 1011010110\,10110 101 discrete parameter combinations (Thomaser et al., 2023).
We construct the surrogate model $\hat{f}^{\mkern 1.0mu\text{S}}$ for the source functions to validate our proposed transfer learning approach on this dataset by utilizing the entire $f^{\text{S}}$ dataset. The complete dataset is also employed as the test set to compute the SMAPE for different GPR models on the target function $f^{\text{T}}$. The transfer dataset $\mathcal{T}$ is randomly sampled from the target function, with a maximum of up to $50$ points. A GPR model is also trained from scratch using only the same transfer dataset without incorporating any prior knowledge. Additionally, we analyze how varying the size of $\mathcal{T}$, ranging from 5 to $50$ points, impacts the performance of the transfer learning approach. Additionally, we include a reproduced baseline method that focuses exclusively on transfer learning using affine transformations only for this dataset (Pan et al., 2024).
@@
We measure the models’ performance using the symmetric mean absolute percentage error (SMAPE) (Flores, 1986). To ensure robustness and reliability, the transfer learning process is conducted ten times for each BBOB function, using randomly generated beta CDF and affine transformation in each iteration. Likewise, the real-world application in the automotive industry is repeated ten times to account for variability.
@@
The BBOB functions are accessed via the IOHexperimenter framework (de Nobel et al., 2024). We implement a GPR model with a Gaussian kernel with the GPy package. [^1] https://gpy.readthedocs.io/en/deploy/ To mitigate the skewness of the function values, a $log$-transformation is applied thereto before training the surrogate model for both synthetic and real-world problems. We use an exponentially decaying learning rate scheduler for the mini-batch gradient descent to minimize the loss function. Hyperparameter tuning for the transfer learning procedure included adjusting the learning rate ($[10^{-3},1]$), batch size ([$0.1|\mathcal{T}|,0.2|\mathcal{T}|$]), number of epochs ($[60,100]$), and the decay rate of the exponential scheduler ($[5\times 10^{-3},0.3]$). These hyperparameters are independently fine-tuned for each BBOB function using the SMAC3 library (Lindauer et al., 2022). The experimental setup details and implementations are available via Zenodo (Authors, s). Additionally, a copy of the supplementary materials is also included.
@@ 5. Experimental Results
Table 1. On 2D BBOB functions, we compare the SMAPE value (mean $\pm$ standard deviation) of three GPR models: the original model, the transferred model, and the one trained from scratch on the transfer dataset. Three transfer sample sizes are investigated $\lvert\mathcal{T}\rvert\in\{20,40,80\}$. The transfer target is created with an exponential-shaped beta CDF. We apply the Kruskal-Wallis test with a significance level of 5%, followed by Dunn’s post-hoc analysis to detect significant winners: the transferred model is underlined if it outperforms the original model; the boldface indicates the better one between the transferred and the model trained from scratch.
@@ 5.1. Transferring GPR on BBOB
Additional analyses in Fig. 8 and Fig. 9 (refer to the supplementary material (Authors, s)) extend these observations to 5D and 10D BBOB functions. In the 5D case, models trained from scratch demonstrate a more noticeable advantage compared to the 2D scenario, surpassing transferred models on certain BBOB functions with as few as 20 samples. This could be because the 5D case involves more complex relations, which the transferred models find harder to capture than the simpler or more closely related 2D case. Interestingly, in the 10D experiments, models trained from scratch exhibit significant performance drops with only 40 samples. A similar issue is observed for some BBOB functions even with 80 samples, as such sample sizes are insufficient to train accurate GPR models (Bull, 2011; Van Der Vaart and Van Zanten, 2008). These results align with prior studies (Pan et al., 2024), which emphasize the value of transfer learning in data-scarce settings but highlight its diminishing returns as data availability grows.
@@
Next, we show the detailed performance values of the 2D scenario in Table 1. Results for 5D and 10D cases are in the supplementary material (Authors, s). To assess the impact of transfer sample size, we compare models using the Kruskal-Wallis test and Dunn’s post-hoc analysis (5% significance). Significant results are highlighted: the transferred model is underlined when it outperforms the original, and the better model between transferred and trained-from-scratch is highlighted in boldface when statistically significant.
@@
For 2D functions, the transferred GPR generally outperforms the model trained from scratch and the original GPR when only 20 transfer samples are available, with F5 and F24 being notable exceptions. However, as the transfer sample size increases to 80, most functions perform better with the model trained from scratch than the transferred model. Despite this, the transferred model consistently delivers significant improvements over the original GPR across all transfer sample sizes, except for F23 and F24. Fig. 4 further explores the relation between model performance and transfer sample sizes. The SMAPE trends show that the transferred model consistently outperforms the model trained from scratch up to 40 sample points for most functions, after which the advantage gradually diminishes. This demonstrates the effectiveness of our transfer learning approach in low-data scenarios, as anticipated. Interestingly, for F5, there is a relatively simple linear function—the model trained from scratch surpasses the transferred model with as few as 20 samples, likely due to the low sample complexity of F5, enabling accurate predictions with minimal data. A detailed analysis reveals that the original GPR model for F23 and F24 suffers from significant underfitting, particularly in the case of F23, which limits the potential benefits of transfer learning. This finding emphasizes that the effectiveness of the transfer learning approach depends on the original GPR model achieving a baseline level of accuracy. Furthermore, the SMAPE trends for F24 highlight specific limitations of the proposed method, indicating that while effective overall, it may struggle with particular functions.
@@
We now analyze how the dimensionality of a function’s domain influences the effectiveness of transfer learning. For 5D functions (refer to Fig. 10 and Table 2 in the supplementary material (Authors, s)), the transferred model significantly outperforms the original GPR across most BBOB functions, regardless of the transfer sample size. However, exceptions are observed in functions F16, F21, and F23. As the transfer data size increases to 40, a growing number of BBOB functions favor models trained from scratch, reducing the relative benefits of transfer learning compared to the 2D scenario. When the transfer sample size reaches 200, the model trained from scratch significantly outperforms the transferred model on several functions. For instance, on F5, the trained-from-scratch model consistently delivers superior performance, regardless of sample size. Furthermore, for functions such as F16 and F23, the performance of both models remains nearly identical across all sample sizes, showing no improvement despite an increase in the number of transfer samples. This is likely because these functions are inherently tricky for GPR to learn effectively.
@@
For 10D functions (refer to the supplementary material (Authors, s)), models trained from scratch struggle with poor performance when sample sizes are small. As dimensionality increases, the complexity of the function landscape requires substantially more data for GPR models to achieve accurate approximations, as previously noted. Under such data-scarce conditions, transfer learning provides a clear advantage over training from scratch. Interestingly, a significant phase transition occurs around 80 samples for the scratch-trained model, marked by a sharp drop in its SMAPE value. This significantly narrows the performance gap with the transferred model. At the largest tested sample size (400), the scratch-trained model outperforms the transferred model across most functions. However, for highly multimodal functions like F16 and F23, transfer learning fails to improve the performance of the original GPR model. In contrast, training the GPR model from scratch with 400 samples achieves superior results on these complex functions. This underscores the limitations of transfer learning in handling highly intricate and multimodal landscapes, where large amounts of task-specific data are essential for optimal performance.

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25 Companion), July 14–18, 2025, Malaga, Spaindoi: 10.1145/3712255.3734291isbn: 979-8-4007-1464-1/2025/07ccs: Computing methodologies Transfer learningccs: Theory of computation Gaussian processesccs: Theory of computation Continuous optimization
@@ 1. Introduction
Surrogate modeling (Forrester et al., 2008; Forrester and Keane, 2009; Bhosekar and Ierapetritou, 2018; Tong et al., 2021) is extensively used to replace expensive simulators for reducing computational costs, for instance in automobile industry (Kiani and Yildiz, 2016; Fang et al., 2017; Qiu et al., 2018; Thomaser et al., 2023). Machine learning models are commonly used as surrogates, e.g., Gaussian process regression (GPR) (Satria Palar et al., 2020; Rajaram et al., 2020; Pourmohamad, 2021) and random forest (Wang and Jin, 2020; Antoniadis et al., 2021). Yet building a new surrogate typically requires many expensive samples, in particular when there are many independent variables. Acquiring this data usually implies running expensive simulations or real-world experiments. Therefore, we wish to avoid the cost of acquiring large data sets to build surrogates on a new problem instance. Transfer learning (Pan and Yang, 2010; Weiss et al., 2016; Zhuang et al., 2021) can be used to tackle this issue: with a tiny transfer data set sampled on a new problem (the target), we can learn to tweak an accurate surrogate trained on an old problem (the source), provided certain symmetry/invariances between problems.
Covariance shift (Sugiyama et al., 2007; Shimodaira, 2000; Pathak et al., 2022) is an important type of symmetry, which says that for a regression task to approximate the source function $f^{\text{S}}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$, a target function $f^{\text{T}}$ can be obtained from $f^{\text{S}}$ by transforming the domain thereof. Namely, there exists a bijection $g\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that $f^{\text{T}}=f^{\text{S}}\circ g$. As for the surrogate modeling, the covariance shift implies the predictive distribution $P(y|\mathbf{x})$ remains unchanged between the source and the target while $P(\mathbf{x})$ differs. Previous studies (Pan et al., 2024, 2025a) have specialized with $g$ being an affine transformation, estimating the parameters by minimizing a loss function on a tiny transfer data evaluated on the target.
@@
Figure 1. For the 2D F7 StepEllipsoid function, we show, from left to right: the contour lines of the source function ($f^{\text{S}}$), the target ($f^{\text{T}}$), the original GPR ($\hat{f}^{\mkern 1.0mu\text{S}}$) trained to approximate $f^{\text{S}}$, the GPR model trained from scratch with 40 samples from $f^{\text{T}}$, and the transferred GPR using the same samples. $f^{\text{T}}$ is created from $f^{\text{S}}$ by transforming the domain thereof with affine warping. We show the transfer effect with 80 data points in the last two subplots.
@@
However, the affine transformation $g$ might be too restrictive to model complex real-world scenarios, e.g., non-linearity is necessary in the automobile industry problems (Pan et al., 2025a). Hence, we propose implementing a non-linear function $g$ with a beta cumulative distribution function (CDF) (Snoek et al., 2014). We showcase the preliminary results of our method in Figure 1, where we train a GPR on function F7 from the BBOB benchmark suite. Comparing the transferred model to the one trained from scratch on the target function (fourth and fifth subplots), we see the transferred model is more accurate and resembles the contour lines of the target function. Our contributions are:
@@ 2. Related Works
Saida and Nishio (2023) investigated GPR surrogate model transfer for structural reliability under uncertainties by augmenting the feature space (Daumé, 2009). Zhang et al. (2023) proposed a novel strategy that employs a geodesic flow kernel and knee point-based manifold learning to refine GPRs using high-quality knee solutions from previous tasks, thereby enriching training data and boosting solution precision. In multi-task settings, Cao et al. (2010) developed the Adaptive Transfer Learning algorithm (AT-GP) using a semi-parametric transfer kernel. The Transfer Bayesian Committee Machine (Tr-BCM) (Da et al., 2019) introduced a scalable transfer learning approach by aggregating predictions from lightweight local experts, relaxing assumptions of uniform similarity between tasks. Recent advances include Papez and Quinn (2022), which proposed a probabilistic predictor for global source-target interactions, and Wei et al. (2023), which developed an interpretable multi-source transfer kernel for improved cross-task performance.
@@
Snoek et al. (2014) proposed warping the independent variables with beta CDF to realize non-stationary kernels, where the shape parameters of the beta distribution are inferred with Bayesian estimation (using log-normal priors). Their posterior predictive distribution is obtained by marginalizing the shape parameters. In contrast, we use the beta CDF to capture the non-linear mapping between source and target domains, fitting its shape parameters directly through loss minimization. Cowen-Rivers et al. (2022) introduced Kumaraswamy input warping as a fast, flexible stand‐in for the beta CDF. Du et al. (2017) presented a Hypothesis Transfer Learning framework linking domains via transformation functions, and Zhu et al. (2023) proposed a nonlinear, data‐agnostic mapping to align marginal distributions.
@@ 3. Learning Affine Warping to transfer domains
where $B(\alpha_{i},\beta_{i})$ is the beta function. The beta CDF parameterization for each dimension has two unknown parameters $\alpha_{i}>0,\beta_{i}>0$. Figure 2 illustrates input warping on a 2D sphere where each axis is warped nonlinearly using its own shape parameters.
Figure 2. 2D input warping: The coordinate system is transformed from left to right by beta CDFs with shape parameter $\alpha=1.0558,\beta=1.9339$ for $x$-axis and $\alpha=0.8655,\beta=1.8148$ for $y$-axis. We show the contour lines of a sphere function on the left and its warped version on the right.
@@ 3.2. Transfer non-differentiable surrogates
We also extend our methodology to non-differentiable models like random forests by tuning parameters with Covariance matrix adaptation evolution strategy (CMA-ES) (Emmerich et al., 2018; Hansen, 2006, 2016). CMA-ES can be applied directly to the search space of the translation parameter $\mathbf{v}$ and the beta CDF parameters $\mathbf{\alpha}$ and $\mathbf{\beta}$, which are Euclidean. However, special treatment is needed for $\mathbf{W}$, which lives in a smooth manifold $\operatorname{SO}\!\left(d\right)$. To solve this issue, we consider the Lie group representation $\mathfrak{so}(d)=\{\mathbf{A}\in\mathbb{R}^{d\times d}\colon\mathbf{A}^{\top}=-\mathbf{A}\}$, which is a flat space (with dimension $d(d-1)/2$), and optimize this representation with CMA-ES. A rotation matrix $\mathbf{W}$ can be recovered from its representation $\mathbf{A}$ with the exponential map, i.e., $\mathbf{W}=\operatorname{Exp}(\mathbf{A})$. For each search point $\mathbf{z}\in\mathbb{R}^{d(d-1)/2}$, we have to transform it into a $d\times d$ antisymmetric matrix to preserve the structure of $\mathfrak{so}(d)$: the components in $\mathbf{z}$ are sequentially assigned to the upper triangular of $\mathbf{A}$ (diagonal entries are zero) row by row. The negative value of the transposition of the upper triangular then fills the lower triangular entries.
Figure 3. On 2D BBOB functions, we compare transferred GPR models with those trained from scratch on the transfer dataset. Each cell displays the percentage difference in average SMAPE (%) for a combination of BBOB functions, sample size, and beta CDF shape. Positive values (shown in red) indicate superior performance of the transferred model.
@@ 4. Experimental Settings
We first evaluate our method on the Black-Box Optimization Benchmarking (BBOB) (Hansen et al., 2009a, 2020, b) suite, which consists of 24 continuous, single-objective problems. The BBOB suite has been widely used as a regression benchmark (Singh and Dwivedi, 2018; Tian et al., 2020; Chen et al., 2022; Yang and Affenzeller, 2023) as it reflects real-world regression difficulties. To create synthetic transfer learning problems out of BBOB, we use the first instance of each BBOB function as the source $f^{\text{S}}$, and construct the target $f^{\text{T}}$ by applying a beta CDF transformation, followed by random rotation and translation transformations, to the base function (so we do not use BBOB’s internal instance transformation methods). The shape parameters are sampled from log-normal distributions, i.e., $\log\alpha_{i}\sim\mathcal{N}(\mu_{i}^{\alpha},\sigma^{\alpha}_{i}),\log\beta_{i}\sim\mathcal{N}(\mu_{i}^{\beta},\sigma_{i}^{\beta})$. Inspired by (Snoek et al., 2014), we vary the priors to induce different warp geometries:
@@
We validate our transfer approach on an ABS‐braking dataset from automotive engineering (Thomaser et al., 2023). The search space consists of two ABS parameters, $x_{1}$ and $x_{2}$, yielding 10 1011010110\,10110 101 discrete settings (Thomaser et al., 2023). We construct the surrogate model $\hat{f}^{\mkern 1.0mu\text{S}}$ for the source functions to validate our proposed transfer learning approach on this dataset by utilizing the entire $f^{\text{S}}$ dataset. The complete dataset is also employed as the test set to compute the SMAPE for different GPR models on the target function $f^{\text{T}}$. The transfer dataset $\mathcal{T}$ is randomly sampled from the target function, with a maximum of up to $50$ points. A GPR model is also trained from scratch using only the same transfer dataset without incorporating any prior knowledge. We also sweep $\mathcal{T}$ from 5 to $50$ to study sample‐size effects, and benchmark against an affine‐only transfer baseline (Pan et al., 2024).
@@
Model accuracy is evaluated via the symmetric mean absolute percentage error (SMAPE) (Flores, 1986). To ensure robustness and reliability, the transfer learning process is conducted ten times per BBOB function, using a randomly generated beta CDF and an affine transformation in each iteration. Likewise, the real-world application in the automotive industry is repeated ten times to account for variability.
@@
The BBOB functions are accessed via the IOHexperimenter framework (de Nobel et al., 2024). We implement a GPR model with a Gaussian kernel with the GPy package. [^1] https://gpy.readthedocs.io/en/deploy/ To mitigate the skewness of the function values, a $log$-transformation is applied before training the surrogate model for both synthetic and real-world problems. We use an exponentially decaying learning rate scheduler for the mini-batch gradient descent to minimize the loss function. In addition, a multi-restart strategy is employed to avoid getting stuck in local minima during the optimization. Hyperparameter tuning for the transfer learning procedure included adjusting the learning rate ($[10^{-3},1]$), batch size ([$0.1|\mathcal{T}|,0.2|\mathcal{T}|$]), number of epochs ($[60,100]$), and the decay rate of the exponential scheduler ($[5\times 10^{-3},0.3]$). These hyperparameters are independently fine-tuned for each BBOB function using the SMAC3 library (Lindauer et al., 2022). The experimental setup details, implementations, and a copy of the supplementary materials are available via Zenodo (Pan et al., 2025b).
@@ 5. Experimental Results
Table 1. On 2D BBOB functions, we compare the SMAPE value (mean $\pm$ standard deviation) of three GPR models: the original model, the transferred model, and the one trained from scratch on the transfer dataset. Three transfer sample sizes are investigated $\lvert\mathcal{T}\rvert\in\{20,40,80\}$. The transfer target is created with an exponential-shaped beta CDF. We apply the Kruskal-Wallis test with a significance level of 5%, followed by Dunn’s post-hoc analysis to detect significant winners: the transferred model is underlined if it outperforms the original; and the better of the transferred and one trained from scratch is shown in bold.
@@ 5.1. Transferring GPR on BBOB
Additional analyses in Fig. 8 and Fig. 9 (refer to the supplementary material (Pan et al., 2025b)) extend these observations to 5D and 10D BBOB functions. In the 5D case, models trained from scratch demonstrate a more noticeable advantage compared to the 2D scenario, surpassing transferred models on certain BBOB functions with as few as 20 samples. This could be because the 5D case involves more complex relations, which the transferred models find harder to capture than the simpler or more closely related 2D case. Interestingly, in the 10D experiments, models trained from scratch exhibit significant performance drops with only 40 samples. A similar issue is observed for some BBOB functions even with 80 samples, as such sample sizes are insufficient to train accurate GPR models (Bull, 2011; Van Der Vaart and Van Zanten, 2008). These results align with prior studies (Pan et al., 2024), which emphasize the value of transfer learning in data-scarce settings but highlight its diminishing returns as data availability grows.
@@
Next, we show the detailed performance values of the 2D scenario in Table 1. Results for 5D and 10D cases are in the supplementary material (Pan et al., 2025b). To assess the impact of transfer sample size, we compare models using the Kruskal-Wallis test and Dunn’s post-hoc analysis (5% significance). Significant results are highlighted: the transferred model is underlined when it outperforms the original, and the better model between transferred and trained-from-scratch is highlighted in boldface when statistically significant.
@@
For 2D functions, the transferred GPR generally outperforms the model trained from scratch and the original GPR when only 20 transfer samples are available, with F5 and F24 being notable exceptions. However, as the transfer sample size increases to 80, most functions perform better with the model trained from scratch than the transferred model. Despite this, the transferred model consistently delivers significant improvements over the original GPR across all transfer sample sizes, except for F23 and F24. Fig. 4 further explores the relation between model performance and transfer sample sizes. The SMAPE trends show that the transferred model consistently outperforms the model trained from scratch up to 40 sample points for most functions, after which the advantage gradually diminishes. This demonstrates the effectiveness of our transfer learning approach in low-data scenarios, as anticipated. Interestingly, for F5, a simple linear function, the model trained from scratch outperforms the transferred model with just 20 samples—likely due to F5’s low sample complexity, which allows accurate learning from minimal data. A detailed analysis reveals that the original GPR model for F23 and F24 suffers from significant underfitting, particularly in the case of F23, which limits the potential benefits of transfer learning. This finding emphasizes that the effectiveness of the transfer learning approach depends on the original GPR model achieving a baseline level of accuracy. Furthermore, the SMAPE trends for F24 highlight specific limitations of the proposed method, indicating that while effective overall, it may struggle with particular functions.
@@
We now analyze how the dimensionality of a function’s domain influences the effectiveness of transfer learning. For 5D functions (refer to Fig. 10 and Table 2 in the supplementary material (Pan et al., 2025b)), the transferred model significantly outperforms the original GPR across most BBOB functions, regardless of the transfer sample size. However, exceptions are observed in functions F16, F21, and F23. As the transfer data size increases to 40, a growing number of BBOB functions favor models trained from scratch, reducing the relative benefits of transfer learning compared to the 2D scenario. When the transfer sample size reaches 200, the model trained from scratch significantly outperforms the transferred model on several functions. For instance, on F5, the trained-from-scratch model consistently delivers superior performance, regardless of sample size. On F16 and F23 both methods remain indistinguishable at all sizes, suggesting these landscapes are simply hard for GPR to capture.
@@
For 10D functions (refer to the supplementary material (Pan et al., 2025b)), models trained from scratch struggle with poor performance when sample sizes are small. As dimensionality increases, the complexity of the function landscape requires substantially more data for GPR models to achieve accurate approximations, as previously noted. Under such data-scarce conditions, transfer learning provides a clear advantage over training from scratch. Interestingly, a significant phase transition occurs around 80 samples for the scratch-trained model, marked by a sharp drop in its SMAPE value. This significantly narrows the performance gap with the transferred model. At the largest tested sample size (400), the scratch-trained model outperforms the transferred model across most functions. However, for highly multimodal functions like F16 and F23, transfer learning fails to improve the performance of the original GPR model. In contrast, training the GPR model from scratch with 400 samples achieves superior results on these complex functions. This underscores the limitations of transfer learning in handling highly intricate and multimodal landscapes, where large amounts of task-specific data are essential for optimal performance. In addition, it is noticeable that the original GPR model still outperforms the model trained from scratch on just 40 or 80 samples, despite never seeing target-instance data. This suggests that the original model’s fit to the 10-dimensional BBOB landscapes, learned from sufficient source data, remains relatively good even without further adaptation.